As you can see in the following​ table, demand for heart transplant surgery at Washington General Hospital has increased steadily in the past few​ years:

Year 1 2 3 4 5 Heart Transplants 45.0 50.0 55.0 54.0 58.0 The director of medical services predicted 6 years ago that demand in year 1 would be 41.0 surgeries.   ​

a) Using exponential smoothing with alpha of 0.60 and the given forecast for year​ 1, the forecasts for years 2 through 6 are ​(round your responses to one decimal ​place)​:

Year 1 2 3 4 5 6 Forecast 41.0 43.40 47.4 51.9 53.2 56.1

For the forecast made using exponential smoothing with alpha ​= 0.60 and the given forecast for year​ 1, MAD​ = ?????? surgeries ​(round your response to one decimal place​).
Answer #1

MAD (mean absolute deviation) for forecasts shows the deviation of forecasted demand from actual demand. This is the mean deviation per period in absolute terms between a number of period forecasts and the corresponding period demand.

MAD using exponential smoothing is calculated as below

MAD(i + 1) = ((i) * ABS(Actual(i) - Forecast(i)) + (1 - ((i)) * MAD(i)

Year At Ft At - Ft Abs (At - Ft)
1 45.0 41.0 4.0 4.0
2 50.0 43.4 6.6 6.6
3 55.0 47.4 7.6 7.6
4 54.0 51.9 2.1 2.1
5 58.0 53.2 4.8 4.8
6 56.1
Total 25.1

MAD = Sum (Abs (At - Ft) ) / n = 25.1 / 5 = 5.02.

If you liked my answer, please give like/ thumbs up. Thank you!